OpenStudy (anonymous):

hey guys...how do you do this? What do the solutions of the graph represent? y=x^2-10x+25

OpenStudy (anonymous):

???

OpenStudy (anonymous):

the x-intercepts are the solutions to 0=x^2-10x+25, sorry.

OpenStudy (anonymous):

Sorry but I still don't understand :)

OpenStudy (anonymous):

If you were to graph the equation y=x^2-10x+25, your graph will touch/intersect the x-axis at the x's that make this statement true: 0=x^2-10x+25.

OpenStudy (anonymous):

Can you solve this equation?: 0=x^2-10x+25

OpenStudy (anonymous):

no, this is where i am stuck. my teacher said something to do with p and q but it makes me more confused :(

OpenStudy (anonymous):

Are p and q the roots?

OpenStudy (anonymous):

or sometimes called zeros?

OpenStudy (anonymous):

I don't know what p and q are :(

OpenStudy (anonymous):

You can actually write the equation: 0 = x^2-10x+25 as 0 = (x-p)(x-q)

OpenStudy (anonymous):

The equation can be written as: 0 = x^2-10x+25 0 = (x-5)(x-5) So your roots (zeros) is 5. Your graph will only touch the x axis at x=5 but not go through it.

OpenStudy (anonymous):

Solutions to any equation always represent points simple as that

OpenStudy (anonymous):

Thanks guys you were very helpful :)